OpenStudy (youngnephew34):

What is the value of x in the proportion 6x+1/7=18x-2/14 A) 0 B) 3 c) 2/3 D) 1/14 PLZZZZZ HELPPPPPP FAN AND MEDAL

OpenStudy (anonymous):

you can get them into a common denominator by multiply the top and bottom of the left by 2

OpenStudy (anonymous):

First you want to find the common denominator... 6x+\frac{1}{7}=18x-\frac{1}{7}

OpenStudy (anonymous):

\(\dfrac{6x+1}{7}=\dfrac{18x-2}{14}=\)

OpenStudy (youngnephew34):

ok

OpenStudy (anonymous):

\(\dfrac{6x+1}{7}=\dfrac{?}{14}\)

OpenStudy (youngnephew34):

so it's 6x1

OpenStudy (anonymous):

no we have to multiply the top and bottom by 2 to get a common denominator 6x+1(2)=12x+2 now we have \(\dfrac{12x+2}{14}=\dfrac{18x-2}{14}\implies 12x+2=18x-2\)

OpenStudy (anonymous):

solve that

OpenStudy (anonymous):

g2g bbl

OpenStudy (youngnephew34):

ok